What is the probability of getting 3 sevens in a row on 3 independent rolls of a pair of fair dice, given that the first two rolls were sevens? Provide your answer as a reduced fraction.
Answer #1

Probability = Favorable Outcomes / Total Outcomes

There are 6 sides on a dice. Therefore Probability of getting a 7 = 1 / 6

Given that we have got a seven in the first 2 rolls. Which means these events have occurred, and therefore the probability of each event = 1 * 1 = 1

Therefore the probability of a 7 coming on the 3rd roll, given that you already have a 7 in the first 2 rolls

P = 1 * 1 * (1 / 6) = 1 / 6